OpenStudy (anonymous):

The domension of prism B are 3 times the dimension of prism A..The volume of prism B is 10368cm3....................WHta is the volume of prism A

OpenStudy (anonymous):

\[(l _{1}/l_{2})^{3} = V _{1}/V_{2}\]

OpenStudy (anonymous):

no

OpenStudy (anonymous):

This is what you should do. \[(1/3)^{3} = x/10368\]\[10368/27\]\[Answer = 384cm^{3}\]

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Your welcome. Try to understand the formulas before you ask another similar question. :D